OpenStudy (anonymous):

derivative of 3x - 5 cos (pi x)^2

OpenStudy (lgbasallote):

you know the derivative of 3x right? no problem there?

OpenStudy (anonymous):

yes , that is 3 but how about the second part how will i am going to perform that

OpenStudy (lgbasallote):

\[-5[\cos (\pi x)]^2 \implies -5\cos^2 (\pi x)\] first of all..you can pull out -5 since it's a constant \[-5( \frac{d}{dx} \cos^2 (\pi x))\] do you know the derivative of cos^2 pi x? or no?

mathslover (mathslover):

Differentiate the sum term by term and factor out constants

mathslover (mathslover):

\[\large{3(\frac{d}{dx}(x))-5\frac{d}{dx}(cos^2(\pi x))}\]

mathslover (mathslover):

Well sorry for inturruption ..@lgbasallote you may continue .. sorry

OpenStudy (anonymous):

why did you move the exponent to cos, is that the rule if in the given cos^2 we will move then.. derivative of cos^2 is -2sin?

OpenStudy (lgbasallote):

well it's not a rule...but it's an option..\[(\cos x)^2 = \cos^2 x\]

OpenStudy (lgbasallote):

i just moved it so it's easier...

OpenStudy (lgbasallote):

anyway...derivative of cos^2 is not -2 sin...you have to use power rule..

OpenStudy (anonymous):

@lgbasallote it will be better I think if you will not move the exponent..

OpenStudy (lgbasallote):

hmm ok..

OpenStudy (anonymous):

I mean to say that it will become easy to use power rule there just like in case of \(x^2\)..

OpenStudy (lgbasallote):

\[[\cos (\pi x)]^2\] first do power rule (which is \(\frac{d}{dx} (x^n) = nx^{n-1}\)) so \[\implies 2[\cos (\pi x)]^{2-1}\] do you get this part so far?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

now...that's not yet finished...according to chain rule...you now have to take the derivative of \[\cos (\pi x)\] now..what is the derivative of this?

OpenStudy (anonymous):

-sin(πx), wait is π a constant?

OpenStudy (lgbasallote):

yes pi is a constant

OpenStudy (lgbasallote):

so you're right...now according to chain rule...you now haveto take the derivative of \[(\pi x)\]

OpenStudy (anonymous):

π?

OpenStudy (lgbasallote):

correct. so now you multiply everything together \[-5[2\cos (\pi x) \times -\sin (\pi x) \times \pi]\] what will this give you?

OpenStudy (anonymous):

10πcos(πx)sin(πx)

OpenStudy (lgbasallote):

right!

OpenStudy (lgbasallote):

you can also do it like this \[\implies 5 \pi [2\sin (\pi x) \cos (\pi x)]\] \[\implies 5\pi \sin (2\pi x)\] because of the trig identity \(2\sin a \cos a = \sin (2a)\)

OpenStudy (anonymous):

is that the final answer? including the 3

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

oh wait , how can we go to 10π^2xsin(πx)^2

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

3+10π^2xsin(πx)^2 that is the answer po. ?

OpenStudy (lgbasallote):

can you draw it? \[10 \pi^2 x \sin (\pi x)^2?\]

OpenStudy (anonymous):

no i cant , that is the answer ksi given on the book?

OpenStudy (lgbasallote):

uhh wait... the qustion is \[\Large 3x - 5\cos (\pi x)^2\] or \[\Large 3x - [5\cos (\pi x)]^2\]

OpenStudy (anonymous):

the first one po

OpenStudy (lgbasallote):

hmm i thought so.... i just solved the second one =_=

OpenStudy (lgbasallote):

okay..here's how to do the first one...

OpenStudy (lgbasallote):

first..pull out -5 then take the derivative of \(\cos (\pi x)^2\)

OpenStudy (lgbasallote):

\[\frac{d}{dx} [\cos (\pi x)^2] \implies -\sin (\pi x)^2 \times 2(\pi x) \times \pi\] do you follow how i got that? or do i need to explain it more?

OpenStudy (anonymous):

I understand it

OpenStudy (lgbasallote):

great.. so \[-5[-\sin(\pi x)^2 \times 2(\pi x) \times pi] \implies - 5[ -\sin (\pi x)^2 \times 2\pi^2 x]\implies 10\pi^2 x \sin (\pi x)^2\] got it?

OpenStudy (lgbasallote):

sorry for screwing it up a while ago =))

OpenStudy (anonymous):

no need for you to say sorry po, I should be the one who apologize, anyways thank you po for helping me. :)

OpenStudy (lgbasallote):

welcome ^_^